OpenStudy (anonymous):

Can someone check my Algebra 2 hw? I give medals!

OpenStudy (anonymous):

x^7+8=28

OpenStudy (anonymous):

I got x=1.53

OpenStudy (anonymous):

\[1.534127404634391^7=20.\]

OpenStudy (anonymous):

yes, thank you! would you mind checking a couple more?

OpenStudy (anonymous):

go ahead.

OpenStudy (anonymous):

sqrt(2x+30) = x+3

OpenStudy (anonymous):

I got x=-21

OpenStudy (anonymous):

Try x = 3

OpenStudy (anonymous):

I did (sqrt(2x+30))^2=(x+3)^2 2x+30=x+3^2 -x -x 1x+30=3^2 x+30=9 -30 -30 x=-21 Can you tell me where I went wrong?

jimthompson5910 (jim_thompson5910):

it looks like you squared (x+3) incorrectly Here is how you should square it (x+3)^2 = (x+3)*(x+3) (x+3)^2 = x*(x+3) + 3*(x+3) (x+3)^2 = x*x + x*3 + 3*x + 3*3 (x+3)^2 = x^2 + 3x + 3x + 9 (x+3)^2 = x^2 + 6x + 9

OpenStudy (anonymous):

okay, thank you!